OpenStudy (anonymous):

Determine whether the graph of the equation is symmetric with respect to the x-axis, the y-axis, and/or the origin. y^2 - x - 1 = 0

OpenStudy (anonymous):

x-axis

OpenStudy (anonymous):

This is symmetric about X-axis as both positive and negative inputs to y are giving same output to x

OpenStudy (ybarrap):

y^2 - x - 1 = 0 y^2 -1 = x x = y^2 - 1 Note that both positive and negative values of y give the same x, visually this means that there are two values of y that give the same x. The value of y for each x will be "reflected" about that value of x, this is true for each vale of x. So this reflection is a property of the symmetry of this equation about the x-axis